Rewrite without using absolute value sign:
|x-2|<4

Found 2 solutions by Edwin McCravy, Theo:
Answer by Edwin McCravy(20054)


   |x-2| < 4

-4 < x-2 < 4
+2    +2  +2
------------
-2 <  x  < 6

Edwin

Answer by Theo(13342)

the basic definition of absolute value is:

abs(x) = x if x is positive and equals -x if x is negative.

x represents any expression that is included as an argument in the absolute value function.

therefore:

abs(x-2) < 4 becomes:

(x-2) < 4 if (x-2) is positive, .....

and -(x-2) < 4 if (x-2) is negative.

when (x-2) is positive, you would solve for x as follows:

x-2 < 4
add 2 to both sides of the inequality to get:
x < 6

when (x-2) is negative, you would solve for x as follows:

-(x-2) < 4
simplify to get:
-x + 2 < 4
subtract 2 from both sides to get:
-x < 2
multiply both sides by -1 to get:
x > -2

don't forget that multiplying both sides of an inequality by the same negative number reverses the inequality.

so your answer is:

x < 6 or x > -2

this can be written as -2 < x < 6

abs(x-2) < 4 when x is greater than 2 and when x is < 6

to test this out, let x equal values that satisfy the solution and that do not satisfy that solution.

any value of x > -2 and < 6 will satisfy the solution.

i choose 4

when x = 4, abs(x-2) < 4 becomes abs(4-2) < 4 which becomes abs(2) < 4 which becomes 2 < 4 which is true.

i choose x = -3 and x = 7 as values which do not satisfy the solution range.

when x = -3, abs(x-2) < 4 becomes abs(-3-2) < 4 which becomes abs(-5) < 4 which becomes 5 < 4 which is false.

when x = 7, abs(x-2) < 4 becomes abs(7-2) < 4 which becomes abs(5) < 4 which becomes 5 < 4 which is false.

you can also graph the equation of y = abs(x) and y = 4 and it will show you graphically when the equation of y = abs(x) is smaller than 4.

that graph is shown below:

$$$

from the graph, you can see that the value of y = abs(x-2) is less than the value of y = 4 when x is greater than -2 through x is smaller than 6 only.

the coordinate points are shown in the format of (x,y), where x represents the value of the x-coordinate and y represents the value of the y coordinate.

for example, when x = -2, then the value of y = abs(x-2) becomes y = abs(-2-2) which becomes y = abs(-4) which becomes y = 4, which provides you with the coordinate point of (-2,4).